# references.bib

### v2

@@ 1 Introduction
Scaling up generalizable robotic imitation learning in the real world is key to developing robust policies for practical applications [position_realworld, pi0]. While teleoperation is widely used for collecting demonstrations, it requires physical robots to record both observations and actions for robot-centric demonstrations, making it costly and difficult to scale. To address this, recent work explores alternative, low-cost data sources such as human videos [gen2act, phantom, avid, instant_policy, mimicplay, atm, learning_by_watching] and in-the-wild demonstrations [umi, airexo_v1, dobbe_stickv1, dexcap, demoat]. These human-centric demonstrations eliminate the need for a robot during collection, reducing costs and improving scalability. Human video approaches typically infer actions via hand or object pose estimation [ditto, r+x, instant_policy, im2flow2act], while in-the-wild demonstrations use devices that interface human motion with robot-relevant data during task execution.
@@
Current in-the-wild demonstration devices fall into two main categories: handheld [umi, rum_stickv2, legato, dobbe_stickv1, demoat] and whole-body devices [arcap, airexo_v1, m2r, dexcap]. Handheld devices are typically designed with end-effectors identical to those of the robot. By equipping both the handheld devices and the robot with in-hand cameras, these methods leverage in-hand observations to ensure visual consistency. However, they face two key issues: (1) reliance on visual SLAM for pose estimation leads to action inaccuracies (§4.3); and (2) limited camera field of view hinders depth perception during interactions (Appendix C.3). On the contrary, whole-body devices offer more accurate action capture and flexible observations, but they typically require fine-tuning with teleoperated data [airexo_v1, dexcap], as in-the-wild demonstrations still exhibit a domain gap compared to real robot data.
@@
To address these challenges, we present the AirExo-2 system for large-scale in-the-wild demonstration collection and adaptation. As illustrated in Fig. LABEL:fig:teaser (left), we introduce demonstration adaptors that align both observations and actions from in-the-wild demonstrations with the robot domain, enabling direct use in imitation learning. On the hardware side, we build upon AirExo [airexo_v1] with a redesigned exoskeleton featuring a stronger mechanical structure and improved calibration, enhancing action precision and system robustness. We conduct comprehensive analyses to evaluate the system’s efficiency and accuracy in demonstration collection.
@@
Beyond scaling up in-the-wild demonstration collection, developing a robust policy is key to fully leveraging these demonstrations. We introduce RISE-2, a generalizable policy that fuses 2D and 3D perception, as illustrated in Fig. LABEL:fig:teaser (right). RISE-2 outperforms prior imitation policies in both in-domain and generalization settings. Notably, it matches the performance of teleoperated-data-trained policies using only in-the-wild demonstrations collected and adapted by AirExo-2 — without any robot data. Under equivalent data collection time, it even surpasses teleoperated baselines. These results highlight the strong potential of combining AirExo-2 and RISE-2 for scalable, generalizable imitation learning. The hardware and codebase will be open-sourced.
@@ 2 Related Works
Demonstration data is fundamental to advancing generalizable imitation learning in robotic manipulation. Recent studies on data scaling laws [data_scaling_law] reveal that imitation learning in robotics follows similar scaling trends observed in natural language processing [scaling_nlp] and computer vision [scaling_video, dit] fields, highlighting the importance of scaling up demonstration collection. Teleoperation is the most direct and widely used approach for collecting real-world robot demonstrations, as it captures data in the robot’s native observation domain and action space. However, scaling up teleoperated demonstration collection is challenging, primarily limited by the high cost and infrastructure demands of deploying more robots. Consequently, constructing large-scale teleoperated robotic datasets [rt1, agibotworld, oxe, rh20t, bcz, droid, bridgedata, robomind] requires substantial human labor and physical resources. Moreover, teleoperation is often less intuitive than natural human manipulation, resulting in higher user learning costs [hajl] and lower demonstration quality [s2i].
To overcome these limitations, recent research explores generating [gensim2, robotwin, gensim, robogen] or augmenting [rocoda, intervengen, dexmimicgen, mimicgen, cyberdemo] demonstrations automatically in simulation, which facilitates demonstration collection but often necessitates sim-to-real adaptation for real-world deployment. Other approaches leverage internet-scale human videos for representation learning [vrb, liv, vip, vc1, r3m, mvp, hrp, mpi] or policy pre-training [gen2act, gr2, vpdd, dexmv, gr1, lapa], yet accurately converting 2D human videos into robot demonstrations with 3D trajectories and robot states remains a challenge [human_video_survey]. Alternatively, in-the-wild data collection uses specialized human-operated devices — such as hand-held grippers [umi, rum_stickv2, legato, dobbe_stickv1, demoat], cameras [ar2d2, eve], VR/AR systems [arcap, egomimic, m2r], motion-capture gloves [dexcap], and exoskeletons [airexo_v1, m2r] — to translate human motions into robot actions.
@@
Despite promising in scalability, the kinematic gap and the visual gap pose obstacles in learning from in-the-wild demonstrations [airexo_v1, m2r]. (1) The kinematic gap refers to discrepancies in motion translation between humans and robots, which can lead to inaccuracies in action execution. This gap can be addressed through visual or mechanical methods. Visual approaches typically use off-the-shelf pose estimation frameworks from commercial devices [arcap, umi, ar2d2, rum_stickv2, egomimic, legato, dobbe_stickv1, dexcap, eve] or structure-from-motion [sfm, demoat] to extract robot poses from demonstrations. In contrast, mechanical approaches [airexo_v1, m2r] employ kinematically isomorphic devices that capture robot poses through angle encoder readings. (2) The visual gap, on the other hand, arises from the presence of specialized devices and human hands in the captured visual data, whereas robot demonstrations must accurately reflect the robot itself. Hand-held devices [umi, rum_stickv2, legato, dobbe_stickv1, demoat] mitigate this by using in-hand cameras and the same end-effector during deployment to maintain visual consistency. Other methods address the visual gap by incorporating robot demonstrations for fine-tuning or co-training [ar2d2, airexo_v1, egomimic, eve], or by employing human-in-the-loop strategies to collect corrective behaviors during policy deployment [arcap, dexcap].
@@
Direct learning from in-the-wild demonstrations highlights the need for a generalizable manipulation policy capable of transferring learned skills to the robot during real-world deployment. A generalizable policy can adapt to new domains or environments, even with limited training data from restricted domains [cage], and should generalize across variations in camera perspectives, backgrounds, objects, and embodiments. While many behavior-cloning-based robotic manipulation policies [bc, dp, act3d, rvt, peract, act] perform well in-domain, they often fail in out-of-distribution scenarios, leading to compounded errors and task failures [rise, cage]. Large-scale pre-training on real-world robot demonstration data can improve generalization [mt-act, rt1, oxe, bcz, openvla, rdt1b, octo, hpt, rt2], but it does not fully address the inherent limit of the policy in adapting across diverse domains. Two promising approaches for improving generalization involve leveraging 3D perceptions [polarnet, act3d, rvt, rise, dp3] to enhance environmental understanding and using 2D foundation models [what_pretrain, spawnnet, soft, cage, same] to enrich the policy’s recognition of complex object and scene information. Recent work [lift3d, sgr, sgrv2] combines these approaches to further enhance policy performance, though challenges remain in effectively integrating these strategies for broader generalization across diverse manipulation domains.
@@ 3.1 AirExo-2: Efficiently Collecting and Adapting In-the-Wild Demonstrations
To this end, we present the AirExo-2 system, designed to efficiently collect and adapt in-the-wild demonstrations for downstream manipulation policy training, as illustrated in Fig. 1. Unlike prior approaches that circumvent the visual gap through in-hand cameras [umi, rum_stickv2, legato, dobbe_stickv1, demoat] or rely on pre-training and fine-tuning paradigm [arcap, ar2d2, airexo_v1, egomimic, dexcap, eve], our system explicitly bridges this gap by converting in-the-wild observations into pseudo-robot views using visual adaptors. Combined with accurate motion tracking enabled by the mechanical design, this semi-automatic adaptation process better aligns demonstration data with robot deployment conditions, allowing policies to directly transfer to real-world robot platform. Please refer to Appendix A for implementation details.
@@
As shown in Fig. 1 (left), we upgrade the AirExo design [airexo_v1] with aluminum profiles to improve structural robustness and motion tracking precision. To facilitate subsequent demonstration adaptations, we design the AirExo-2 to have a 1:1 scale and maintain kinematic isomorphism with the robotic arm. A mobile base is added to support data collection in diverse environments, and a dedicated calibration process is introduced to ensure accurate motion capture. Details of the hardware design and calibration procedure are provided in Appendix A.1 and Appendix A.2, respectively. The system retains the low-cost advantage of the original AirExo, with the dual-arm platform (excluding the camera) priced at just \$600.
@@
Following [rise, idp3], we project all states and actions into the global camera coordinate — a universal coordinate frame between the AirExo-2 and robot platform.
@@
As shown in Fig. 1 (top), we convert in-the-wild image into pseudo-robot image by (1) inpainting human-related regions — including hands and AirExo-2 — and (2) overlaying it with the generated robot image. Leveraging the joint encoder readings, calibration results, and the joint mapping between AirExo-2 and the robot, we render RGB-D and mask images of both systems in Open3D [open3d]. The rendered AirExo-2 mask is merged with the hand mask obtained via SAM-2 [sam2] to localize human-related areas, which are then inpainted using ProPainter [propainter]. Inspired by [rovi_aug], the rendered robot image is refined via ControlNet-guided Stable Diffusion [stable_diffusion, controlnet] and composited onto the inpainted image to generate the final photo-realistic pseudo-robot image.
@@ 3.2 RISE-2: A Generalizable Policy for Learning from In-the-Wild Demonstrations
3D point cloud data contains rich spatial structure, enabling effective extraction of local geometric features — a property widely used in general grasping [anygrasp, fang2020graspnet, wang2021graspness]. RISE [rise] encodes point clouds using raw color information to capture both semantic and geometric cues, but it struggles to disentangle coordinate shifts from color variations. RISE-2 inherits the sparse 3D encoder from RISE, but removes color input to focus purely on geometric features. Formally, the sparse encoder $\mathrm{E}_{s}:(D,K)\rightarrow(\mathbf{F}_{g},\mathbf{C}_{g})$, maps a depth image $D$ and camera intrinsics $K$ to sparse geometric features $\mathbf{F}_{g}$, along with the corresponding seed point coordinates $\mathbf{C}_{g}=\{c_{g}^{i}\in\mathbf{P}\}$ after network down-sampling. The encoder adopts a ResNet-like architecture [resnet] built on MinkowskiEngine [minkowski], enabling efficient sparse feature extraction and supporting real-time policy execution.
@@
A generalizable policy requires rich semantic features to understand the scene, while the low-quality texture of point cloud data poses a challenge to this demand. Hence, RISE-2 adopts a dense 2D encoder to focus on semantic understanding based on the color image. Formally, the dense encoder $\mathrm{E}_{d}:(I,D,K)\rightarrow(\mathbf{F}_{s},\mathbf{C}_{s})$ processes the color image $I$ into semantic feature map $\mathbf{F}_{s}=\{f_{s}^{i}\}$ densely organized in the 2D form of shape $h\times w$, along with its reference 3D coordinates $\mathbf{C}_{s}=\{c_{s}^{i}\}$, calculated by applying 2D adaptive average pooling to size $h\times w$ on the point cloud generated from the depth map $D$ and camera intrinsics $K$. One key advantage of the dense encoder is its compatibility with 2D foundation models, which provides highly generalizable visual representations across diverse tasks and domains [cage]. We implement $\mathrm{E}_{d}$ using DINOv2 [dino] fine-tuned with LoRA [lora], enhancing the robustness and adaptability of the policy in capturing contextual relationships within the environment.
@@
RISE-2 extracts geometric and semantic features using separate encoders, posing a challenge for feature fusion from different domains. A straightforward solution is to concatenate the aggregated feature vectors, but this discards fine-grained local information critical for precise perception. An alternative [sgr, sgrv2] upsamples $\mathbf{F}_{s}$ to match the resolution of the input image $I$, projects it onto the point cloud, and then downsamples to align with the seed points $\mathbf{C}_{g}$. However, this process is computationally expensive and reduces the efficiency of policy training and deployment. Instead, RISE-2 utilizes a spatial aligner to efficiently fuse two kinds of features based on their 3D coordinates $\mathbf{C}_{g}$ and $\mathbf{C}_{s}$. For each point $c_{g}^{i}\in\mathbf{C}_{g}$ output by the sparse encoder $\mathrm{E}_{s}$, we compute its $M$ nearest neighbors $\mathbf{N}_{i}=\{n^{i}_{1},n^{i}_{2},\cdots,n^{i}_{M}\}\subseteq\mathbf{C}_{s}$ from the coordinates output by the dense encoder $\mathrm{E}_{d}$. The aligned semantic feature $f^{i}_{s*}$ of $c_{g}^{i}$ is computed by weighted spatial interpolation:
@@
where $\mathrm{dist}(c_{i},c_{j})$ denotes the Euclidean distance between $c_{i}$ and $c_{j}$. Then, the aligned feature for point $c_{s}^{i}$ is obtained by concatenation, i.e., $f_{i}=\mathrm{Concat}(f_{g}^{i},f_{s*}^{i})$. By aligning the seed points $\mathbf{C}_{g}$ with the dense semantic feature map $\mathbf{F}_{s}$ through their 3D coordinates, we can accurately retrieve fine-grained semantic features at arbitrary locations. The fused features are then aggregated into high-level sparse representations using sparse convolution layers [minkowski]. Visualization of the weighted spatial interpolation applied to the 2D feature map is provided in Appendix B.3.
@@
RISE-2 utilizes a decoder-only transformer [vaswani2017attention] to map point tokens with sparse positional encoding [rise] to the latent feature, which conditions a diffusion head [dp, ddpm, janner2022planning] to generate the action chunk [act]. Actions are represented in the camera frame for cross-scene consistency, with translations in absolute coordinates and rotations in the 6D format [rot6d].

### v3

# AirExo-2: Scaling up Generalizable Robotic Imitation Learning with Low-Cost Exoskeletons

@@ 1 Introduction
Scaling up generalizable robotic imitation learning in the real world is key to developing robust policies for practical applications [4, 7]. While teleoperation is widely used for collecting demonstrations, it requires physical robots to record both observations and actions for robot-centric demonstrations, making it costly and difficult to scale. To address this, recent work explores alternative, low-cost data sources such as human videos [5, 57, 91, 97, 99, 108, 112] and in-the-wild demonstrations [19, 29, 88, 100, 115]. These human-centric demonstrations eliminate the need for a robot during collection, reducing costs and improving scalability. Human video approaches typically infer actions via hand or object pose estimation [36, 76, 97, 113], while in-the-wild demonstrations use devices that interface human motion with robot-relevant data during task execution.
@@
Current in-the-wild demonstration devices fall into two main categories: handheld [19, 24, 87, 88, 115] and whole-body devices [17, 29, 53, 100]. Handheld devices are typically designed with end-effectors identical to those of the robot. By equipping both the handheld devices and the robot with in-hand cameras, these methods leverage in-hand observations to ensure visual consistency. However, they face two key issues: (1) reliance on visual SLAM for pose estimation leads to action inaccuracies (§4.3); and (2) limited camera field of view hinders depth perception during interactions (Appendix C.3). On the contrary, whole-body devices offer more accurate action capture and flexible observations, but they typically require fine-tuning with teleoperated data [29, 100], as in-the-wild demonstrations still exhibit a domain gap compared to real robot data.
@@
To address these challenges, we present the AirExo-2 system for large-scale in-the-wild demonstration collection and adaptation. As illustrated in Fig. LABEL:fig:teaser (left), we introduce demonstration adaptors that align both observations and actions from in-the-wild demonstrations with the robot domain, enabling direct use in imitation learning. On the hardware side, we build upon AirExo [29] with a redesigned exoskeleton featuring a stronger mechanical structure and improved calibration, enhancing action precision and system robustness. We conduct comprehensive analyses to evaluate the system’s efficiency and accuracy in demonstration collection.
@@
Beyond scaling up in-the-wild demonstration collection, developing a robust policy is key to fully leveraging these demonstrations. We introduce RISE-2, a generalizable policy that fuses 2D and 3D perception, as illustrated in Fig. LABEL:fig:teaser (right). RISE-2 outperforms prior imitation policies in both in-domain and generalization settings. Notably, it matches the performance of teleoperated-data-trained policies using only in-the-wild demonstrations collected and adapted by AirExo-2 — without any robot data. Under equivalent data collection time, it even surpasses teleoperated baselines. These results highlight the strong potential of combining AirExo-2 and RISE-2 for scalable, generalizable imitation learning. The hardware and codebase are open-sourced at the project website.
@@ 2 Related Works
Demonstration data is fundamental to advancing generalizable imitation learning in robotic manipulation. Recent studies on data scaling laws [59] reveal that imitation learning in robotics follows similar scaling trends observed in natural language processing [49] and computer vision [35, 77] fields, highlighting the importance of scaling up demonstration collection. Teleoperation is the most direct and widely used approach for collecting real-world robot demonstrations, as it captures data in the robot’s native observation domain and action space. However, scaling up teleoperated demonstration collection is challenging, primarily limited by the high cost and infrastructure demands of deploying more robots. Consequently, constructing large-scale teleoperated robotic datasets [9, 10, 21, 27, 45, 52, 98, 110] requires substantial human labor and physical resources. Moreover, teleoperation is often less intuitive than natural human manipulation, resulting in higher user learning costs [65] and lower demonstration quality [13].
To overcome these limitations, recent research explores generating [43, 72, 105, 107] or augmenting [1, 40, 48, 70, 103] demonstrations automatically in simulation, which facilitates demonstration collection but often necessitates sim-to-real adaptation for real-world deployment. Other approaches leverage internet-scale human videos for representation learning [2, 67, 68, 69, 73, 81, 93, 118] or policy pre-training [5, 12, 32, 80, 109, 114], yet accurately converting 2D human videos into robot demonstrations with 3D trajectories and robot states remains a challenge [71]. Alternatively, in-the-wild data collection uses specialized human-operated devices — such as hand-held grippers [19, 24, 87, 88, 115], cameras [23, 104], VR/AR systems [17, 50, 53], motion-capture gloves [100], and exoskeletons [29, 53] — to translate human motions into robot actions.
@@
Despite promising in scalability, the kinematic gap and the visual gap pose obstacles in learning from in-the-wild demonstrations [29, 53]. (1) The kinematic gap refers to discrepancies in motion translation between humans and robots, which can lead to inaccuracies in action execution. This gap can be addressed through visual or mechanical methods. Visual approaches typically use off-the-shelf pose estimation frameworks from commercial devices [17, 19, 23, 24, 50, 87, 88, 100, 104] or structure-from-motion [86, 115] to extract robot poses from demonstrations. In contrast, mechanical approaches [29, 53] employ kinematically isomorphic devices that capture robot poses through angle encoder readings. (2) The visual gap, on the other hand, arises from the presence of specialized devices and human hands in the captured visual data, whereas robot demonstrations must accurately reflect the robot itself. Hand-held devices [19, 24, 87, 88, 115] mitigate this by using in-hand cameras and the same end-effector during deployment to maintain visual consistency. Other methods address the visual gap by incorporating robot demonstrations for fine-tuning or co-training [23, 29, 50, 104], or by employing human-in-the-loop strategies to collect corrective behaviors during policy deployment [17, 100].
@@
Direct learning from in-the-wild demonstrations highlights the need for a generalizable manipulation policy capable of transferring learned skills to the robot during real-world deployment. A generalizable policy can adapt to new domains or environments, even with limited training data from restricted domains [111], and should generalize across variations in camera perspectives, backgrounds, objects, and embodiments. While many behavior-cloning-based robotic manipulation policies [78, 18, 30, 31, 90, 123] perform well in-domain, they often fail in out-of-distribution scenarios, leading to compounded errors and task failures [102, 111]. Large-scale pre-training on real-world robot demonstration data can improve generalization [6, 9, 21, 45, 55, 62, 74, 106, 127], but it does not fully address the inherent limit of the policy in adapting across diverse domains. Two promising approaches for improving generalization involve leveraging 3D perceptions [16, 30, 31, 102, 116] to enhance environmental understanding and using 2D foundation models [11, 60, 79, 111, 119] to enrich the policy’s recognition of complex object and scene information. Recent work [47, 121, 122] combines these approaches to further enhance policy performance, though challenges remain in effectively integrating these strategies for broader generalization across diverse manipulation domains.
@@ 3.1 AirExo-2: Efficiently Collecting and Adapting In-the-Wild Demonstrations
To this end, we present the AirExo-2 system, designed to efficiently collect and adapt in-the-wild demonstrations for downstream manipulation policy training, as illustrated in Fig. 1. Unlike prior approaches that circumvent the visual gap through in-hand cameras [19, 24, 87, 88, 115] or rely on pre-training and fine-tuning paradigm [17, 23, 29, 50, 100, 104], our system explicitly bridges this gap by converting in-the-wild observations into pseudo-robot views using visual adaptors. Combined with accurate motion tracking enabled by the mechanical design, this semi-automatic adaptation process better aligns demonstrations with robot deployment conditions, allowing policies to directly transfer to real-world robot platform. Please refer to Appendix A for implementation details.
@@
As shown in Fig. 1 (left), we upgrade the AirExo design [29] with aluminum profiles to improve structural robustness and motion tracking precision. To facilitate subsequent demonstration adaptations, we design the AirExo-2 to have a 1:1 scale and maintain kinematic isomorphism with the robotic arm. A mobile base is added to support data collection in diverse environments, and a dedicated calibration process is introduced to ensure accurate motion capture. Details of the hardware design and calibration procedure are provided in Appendix A.1 and Appendix A.2, respectively. The system retains the low-cost advantage of the original AirExo, with the dual-arm platform (excluding the camera) priced at just \$600.
@@
Following [102, 117], we project all states and actions into the global camera coordinate — a universal coordinate frame between the AirExo-2 and robot platform.
@@
As shown in Fig. 1 (top), we convert in-the-wild image into pseudo-robot image by (1) inpainting human-related regions — including hands and AirExo-2 — and (2) overlaying it with the generated robot image. Leveraging the joint encoder readings, calibration results, and the joint mapping between AirExo-2 and the robot, we render RGB-D and mask images of both systems in Open3D [124]. The rendered AirExo-2 mask is merged with the hand mask obtained via SAM-2 [82] to localize human-related areas, which are then inpainted using ProPainter [125]. Inspired by [14], the rendered robot image is refined via ControlNet-guided Stable Diffusion [83, 120] and composited onto the inpainted image to generate the final photo-realistic pseudo-robot image.
@@ 3.2 RISE-2: A Generalizable Policy for Learning from In-the-Wild Demonstrations
3D point cloud data contains rich spatial structure, enabling effective extraction of local geometric features — a property widely used in general grasping [26, 25, 101]. RISE [102] encodes point clouds using raw color information to capture both semantic and geometric cues, but it struggles to disentangle coordinate shifts from color variations. RISE-2 inherits the sparse 3D encoder from RISE, but removes color input to focus purely on geometric features. Formally, the sparse encoder $\mathrm{E}_{s}:(D,K)\rightarrow(\mathbf{F}_{g},\mathbf{C}_{g})$, maps a depth image $D$ and camera intrinsics $K$ to sparse geometric features $\mathbf{F}_{g}$, along with the corresponding seed point coordinates $\mathbf{C}_{g}=\{c_{g}^{i}\in\mathbf{P}\}$ after network down-sampling. The encoder adopts a ResNet-like architecture [33] built on MinkowskiEngine [20], enabling efficient sparse feature extraction and supporting real-time policy execution.
@@
A generalizable policy requires rich semantic features to understand the scene, while the low-quality texture of point cloud data poses a challenge to this demand. Hence, RISE-2 adopts a dense 2D encoder to focus on semantic understanding based on the color image. Formally, the dense encoder $\mathrm{E}_{d}:(I,D,K)\rightarrow(\mathbf{F}_{s},\mathbf{C}_{s})$ processes the color image $I$ into semantic feature map $\mathbf{F}_{s}=\{f_{s}^{i}\}$ densely organized in the 2D form of shape $h\times w$, along with its reference 3D coordinates $\mathbf{C}_{s}=\{c_{s}^{i}\}$, calculated by applying 2D adaptive average pooling to size $h\times w$ on the point cloud generated from the depth map $D$ and camera intrinsics $K$. One key advantage of the dense encoder is its compatibility with 2D foundation models, which provides highly generalizable visual representations across diverse tasks and domains [111]. We implement $\mathrm{E}_{d}$ using DINOv2 [75] fine-tuned with LoRA [42], enhancing the robustness and adaptability of the policy in capturing contextual relationships within the environment.
@@
RISE-2 extracts geometric and semantic features using separate encoders, posing a challenge for feature fusion from different domains. A straightforward solution is to concatenate the aggregated feature vectors, but this discards fine-grained local information critical for precise perception. An alternative [121, 122] upsamples $\mathbf{F}_{s}$ to match the resolution of the input image $I$, projects it onto the point cloud, and then downsamples to align with the seed points $\mathbf{C}_{g}$. However, this process is computationally expensive and reduces the efficiency of policy training and deployment. Instead, RISE-2 utilizes a spatial aligner to efficiently fuse two kinds of features based on their 3D coordinates $\mathbf{C}_{g}$ and $\mathbf{C}_{s}$. For each point $c_{g}^{i}\in\mathbf{C}_{g}$ output by the sparse encoder $\mathrm{E}_{s}$, we compute its $M$ nearest neighbors $\mathbf{N}_{i}=\{n^{i}_{1},n^{i}_{2},\cdots,n^{i}_{M}\}\subseteq\mathbf{C}_{s}$ from the coordinates output by the dense encoder $\mathrm{E}_{d}$. The aligned semantic feature $f^{i}_{s*}$ of $c_{g}^{i}$ is computed by weighted spatial interpolation:
@@
where $\mathrm{dist}(c_{i},c_{j})$ denotes the Euclidean distance between $c_{i}$ and $c_{j}$. Then, the aligned feature for point $c_{s}^{i}$ is obtained by concatenation, i.e., $f_{i}=\mathrm{Concat}(f_{g}^{i},f_{s*}^{i})$. By aligning the seed points $\mathbf{C}_{g}$ with the dense semantic feature map $\mathbf{F}_{s}$ through their 3D coordinates, we can accurately retrieve fine-grained semantic features at arbitrary locations. The fused features are then aggregated into high-level sparse representations using sparse convolution layers [20]. Visualization of the weighted spatial interpolation applied to the 2D feature map is provided in Appendix B.3.
@@
RISE-2 utilizes a decoder-only transformer [96] to map point tokens with sparse positional encoding [102] to the latent feature, which conditions a diffusion head [18, 37, 46] to generate the action chunk [123]. Actions are represented in the camera frame for cross-scene consistency, with translations in absolute coordinates and rotations in the 6D format [126].